OpenStudy (anonymous):

The volume in cubic feet of a workshop's storage chest can be expressed as the product of it's three dimensions: V(x) = x^3 - 3x^2 - x +3. The depth is x + 1 Help find a.) Find linear expression with integer coefficients for the other dimensions. b.) If the depth of the chest is 6 feet, what are the other dimensions. Please Help me and show your work.

OpenStudy (anonymous):

oh you know synthetic divison

OpenStudy (anonymous):

if you have one of the roots you can divide into the polynomial to get a quadratic and solve that

OpenStudy (anonymous):

No, I really dont know how to do synthetic divison :/

OpenStudy (anonymous):

Oh ok.

OpenStudy (anonymous):

well since x + 1 works we should try putting -1 in

OpenStudy (anonymous):

\[(-1)^{3} - 3(-1)^{2} - (-1) + 3 = 0\]

OpenStudy (anonymous):

c0=f(-1)=0 c1=f'(-1)=8 f_L(x)=8(x+1)

OpenStudy (anonymous):

Which is that for a or b ?

OpenStudy (anonymous):

a+b(x+1) so a=0 and b=8

OpenStudy (anonymous):

or ax+b a=8 and b=8

OpenStudy (anonymous):

I mean there is two questions, question a and b. I needed both of them explained to me.

OpenStudy (anonymous):

on how to find them